OpenStudy (anonymous):

a=-2 , b=-1 , c=4 , d=3 Solve)c2-b

OpenStudy (callisto):

\[c^2 -b\] Is this your question?

OpenStudy (anonymous):

yes

OpenStudy (callisto):

Put c=4 and b=-1 into the... statement? \[c^2 -b = (4)^2 - (-1) = 16 +1 =?\]